OpenStudy (anonymous):

What is f(-3) if f(x)=|2x+1|+5

OpenStudy (jdoe0001):

\(\bf f(\color{red}{x})=|2\color{red}{x}+1|+5\\ \quad \\ f(\color{red}{-3}) = |2\color{red}{(-3)}+1|+5\)

OpenStudy (anonymous):

ok so then what do i do next? cause i get confused with these problems

OpenStudy (jdoe0001):

well, just multiply the 2 times -3, and expand keep in mind that whatever you get inside the | bars |, will come out POSITIVE anyway absolute values are always positive

OpenStudy (anonymous):

so i do 2*-3 to get -6 then add the 1. but what do i do with the -3 and the f

OpenStudy (jdoe0001):

f(-3) just means whenever you encounter the "x" in the function, replace it for -3 which we already did

OpenStudy (anonymous):

oh ok thanks for the help

OpenStudy (jdoe0001):

\(\bf f(x)=|2x+1|+5\\ \quad \\ f(-3) = |2(-3)+1|+5\implies f(-3) = |-6+1|+5\\\quad \\ \implies f(-3) = |-5|+5\implies f(-3) = 5+5\)

OpenStudy (anonymous):

thank you.

OpenStudy (jdoe0001):

yw